Grade 12th passGeneral Physics

the least count of stop watch is 0.1 second..the time for 20 vibration is 25 second..the maximum uncertainty in time period will be?

sadia
8 Years agoGrade 12th pass

1 Answer

Khimraj
8 Years ago
total time = 25 s ± 0.1 s and error in one oscillation = ( 25 s ± 0.1 s)/ 20 = 1.25 s ± 0.005 s Then percentage error = (0.005/1.25)*100 = 0.4% Hope it clears.